How much acid must be added to 75 grams of solution that is 65% acid to obtain a new solution that is 75% acid?
Answer by CharlesG2(834)

How much acid must be added to 75 grams of solution that is 65% acid to obtain a new solution that is 75% acid?
% in A * amount A + % in B * amount B = % in C * amount C
100% * x + 65% * 75 = 75% * (x + 75)
x + (0.65)(75) = 0.75x + (0.75)(75)
x + 48.75 = 0.75x + 56.25
0.25x = 7.5
x = 30 grams